need help on this
solve for x by completing the square;
X^2-6x+13=0

Answer by jai_kos(139)

x^2 -6x +13 = 0
This can be written as
(x^2 – 6x + 9)+ 4 = 0
(x-3)^2 + 4 =0
(x-3)^2 = - 4
Take the square root on both sides, we get
(x-3) = sqrt(-4), we get
x = sqrt(-4) + 3